# SPARC: Prediction-Based Safe Control for Coupled Controllable and Uncontrollable Agents with Conformal Predictions

### v1

@@ Abstract
We investigate the problem of safe control synthesis for systems operating in environments with uncontrollable agents whose dynamics are unknown but coupled with those of the controlled system. This scenario naturally arises in various applications, such as autonomous driving and human-robot collaboration, where the behavior of uncontrollable agents, like pedestrians, cannot be directly controlled but is influenced by the actions of the autonomous vehicle or robot. In this paper, we present SPARC (Safe Prediction-Based Robust Controller for Coupled Agents), a novel framework designed to ensure safe control in the presence of coupled uncontrollable agents. SPARC leverages conformal prediction to quantify uncertainty in data-driven prediction of agent behavior. Particularly, we introduce a joint distribution-based approach to account for the coupled dynamics of the controlled system and uncontrollable agents. By integrating the control barrier function (CBF) technique, SPARC provides provable safety guarantees at a high confidence level. We illustrate our framework with a case study involving an autonomous driving scenario with walking pedestrians.
@@ I INTRODUCTION
Autonomous systems are increasingly being deployed in our society with a wide range of applications, from self-driving cars to industrial robots. Ensuring safety is one of the most critical tasks in these systems, as they must operate in dynamic environments while avoiding collisions and other unsafe behaviors. For example, in autonomous driving, the system must continuously assess its surroundings to ensure the safety of pedestrians, other vehicles, and passengers.
To achieve safe control in real-world scenarios, one must take the presence of uncontrollable agents in the environment into account. These agents, such as pedestrians or other vehicles, move and make decisions independently of the autonomous system. For instance, in an urban driving scenario, a self-driving car must account for the unpredictable behavior or intention of pedestrians, who may change speed or direction in response to the car’s actions. Managing this interaction is crucial to ensuring the overall safety of the system.
There are two major challenges in achieving safe control in the presence of uncontrollable agents. First, the dynamics of the controlled system and the uncontrollable agents are generally coupled. In other words, the behavior or intention of the uncontrollable agent depends on the state of the controlled system. For example, in autonomous driving, a pedestrian’s movements may change in response to the position and speed of a nearby vehicle. Second, the intentions of such uncontrollable agent are often unknown a priori or very difficult to model precisely. As a result, traditional model-based approaches, such as robust control, cannot be reliably used to ensure safety. The inability to predict the agent’s exact behavior makes it challenging to design a control strategy that accounts for all possible interactions with the uncontrollable agent.
Data-driven approaches provide a powerful tool for handling the unknown intentions of uncontrollable agents. By leveraging large datasets collected from real-world interactions, one can develop highly accurate predictors to forecast the trajectories of uncontrollable agents for control purposes. For instance, as shown in Figure 1, in autonomous driving within pedestrian-rich environments, there are extensive datasets that capture how pedestrians react to vehicles under various conditions. However, a common drawback of data-driven approaches is the lack of formal guarantees. While the prediction of a pedestrian’s intention may be correct 95% of the time, the remaining 5% of failures can lead to unacceptable risks. In safety-critical systems, such errors are intolerable, and one has to design a robust system by taking the uncertainty in data-driven approaches into account.
(a) Intersections Drone Dataset (inD) [1]
(b) Stanford Drone Dataset (SDD) [2]
@@
In this paper, we propose a new framework, called SPARC (Safe Prediction-Based Robust Controller for Coupled Agents), for ensuring safe control in the presence of uncontrollable agents. Specifically, we assume the dynamics of the uncontrollable agent are unknown and coupled with the dynamics of the controlled system. To address this challenge, we leverage the theory of conformal prediction to quantify uncertainty in data-driven predictions of the uncontrollable agent’s intentions. Unlike previous approaches, which assume that the dynamics of the uncontrollable agent are independent of the controlled system, we develop a novel method to sample data and construct conformal regions based on the joint distribution over an augmented state space. This approach allows us to capture the inter-dependencies between the two agents and improve prediction accuracy. Furthermore, by incorporating the control barrier function (CBF) technique, we can provide provable safety guarantees with a confidence bound, ensuring robust system performance. Our framework is particularly well-suited for applications such as autonomous driving in urban environments or human-robot collaborations, where the uncontrollable environments have their own underlying decision-laws that depend on the behavior of the controlled systems.
@@ II Related Works
Safety-Critical Control using CBFs. Control Barrier Functions (CBFs) are a widely used approach for ensuring safety in control systems [3, 4, 5, 6, 7]. First introduced by Ames et al. [8], the key idea behind CBFs is to encode safety constraints into a function that guarantees the system remains within a safe set during operation. However, a significant challenge with CBFs is their reliance on an accurate dynamic model of the system. When the system includes an uncontrollable component with unknown dynamics, applying traditional CBFs can be problematic. In such cases, robust CBFs [9, 10] can be used to account for the worst-case scenario, which, however, often results in overly conservative control strategies.
@@
Conformal Prediction in Control Synthesis. Conformal prediction is an emerging technique in machine learning that provides a way to quantify the uncertainty of model predictions [18]. It works by generating prediction sets or confidence regions that, with a specified level of certainty, are guaranteed to contain the true outcome. This technique is model-free, meaning it can be applied to any underlying predictive model, and it offers formal statistical guarantees on the accuracy of its predictions. Recently, conformal prediction has been extensively applied to control synthesis problems; see, e.g., [19, 20, 21, 22, 23, 24]. Our work is closely related to [22] and [19], which address the problem of safe planning in unknown dynamic environments. In these works, the authors also employ conformal prediction to estimate the confidence region of the uncontrollable agent. However, a key distinction is that these approaches assume the behavior of uncontrollable agents is independent of the controlled system, an assumption that may not hold in many real-world scenarios. In contrast, our framework considers a coupled setting, addressing the issue of distribution shift in conformal prediction, where the behavior of uncontrollable agents is influenced by the state of the controlled system.
@@ III Problem Formulation
Notations: We denote by $\mathbb{R},\mathbb{N}$, and $\mathbb{R}^{n}$ the set of real numbers, natural numbers, and real vectors, respectively. Let $\beta:\mathbb{R}\rightarrow\mathbb{R}$ denote an extended class $\mathcal{K}_{\infty}$ function, i.e., a strictly increasing function with $\beta(0)=0$. For a vector $v\in\mathbb{R}^{n}$, let $\|v\|$ denote its Euclidean norm. For a bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, let $\partial\mathcal{S}$ denote its boundary.
@@ III-A System Models
where $x(t)\in\mathbb{R}^{n}$ represents the state of the system at time $t$, and $u(t)\in\mathcal{U}$ denotes the control input, with $\mathcal{U}\subseteq\mathbb{R}^{m}$ being the set of admissible control inputs. The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ represent the internal system dynamics and the input influence, respectively, and both are assumed to be locally Lipschitz continuous. For convenience, we denote $\mathbf{X}(t)=[x(t),\dot{x}(t)]\in\mathbb{R}^{n_{X}}$.
Uncontrollable Agent: We consider the scenario where the controlled system in Eq. (1) operates among some uncontrollable agents. For the sake of simplicity, we only consider a single uncontrollable agent, and our approach can be extended to $N$ uncontrollable agents by considering their states as a vector. We denote the state of the agent at time instant $t$ by $Y(t)$, where $Y(t)\in\mathbb{R}^{n_{Y}}$. Note that $n_{Y}$ does not necessarily equal $n$, since the dynamics of $X$ and $Y$ may differ.
More specifically, we assume that the uncontrollable agent has its own dynamic and decision law (intention), which can be either deterministic or stochastic. However, we do not assume to have any a priori knowledge of the agent dynamics or intentions. Note that, in general, the behavior of the uncontrollable agent depends on the state of the controlled agent. For example, in autonomous driving scenarios with moving pedestrians (uncontrollable agents), the behavior of the pedestrians is significantly influenced by the position and velocity of the car. Specifically, the closer the car is, the faster pedestrians tend to move away, or the faster the car approaches, the more cautious pedestrians become in their walking strategies.
The objective is to synthesize a control input $u(t)$ for each time instant such that the controlled system can ensure safety with respect to the uncontrollable agent, with a certain level of confidence. Formally, safety here can be framed as a safe set $\mathcal{S}\subseteq\mathbb{R}^{n}\times\mathbb{R}^{n_{Y}}$. Then the problem formulation is as follows.
@@
Consider a controlled system described by Eq. (1) and an uncontrollable agent whose behavior depends on the behavior of the controlled system. Suppose that the controller can access the states of both the controlled system and uncontrollable agent precisely without measurement error, but the physical dynamics and intentions of the uncontrollable agent is unknown a priori. Design a controller that issues control input $u(t)$ for each time instant such that the closed-loop behavior is safe w.r.t. safety set $\mathcal{S}$, with a certain level of confidence.
Since the behavior of the uncontrollable agent is complete unknown, hereafter, we will use sample data to provide confidence level guarantee for safety of form: “the probability that the system becomes unsafe within a certain time horizon is smaller than a threshold bound $\alpha$”.
@@ IV-A Overall Control Framework
  • At each discrete time step, referred to as a prediction update point, we use a neural network to predict the intention of the uncontrollable agent over the next control period $\Delta t$.
  • We apply the technique of conformal predictions to enlarge the predicted results into a region that contains the agents’ reachable states with high confidence.
@@ IV-B Data Collection and Predictor Design
To collect data and to predict the behavior of the uncontrollable agent, we sample the trajectories of both the controlled system and the uncontrollable agent with time interval $\Delta t$. We denote by $[X_{0},X_{1},\ldots,X_{k}]$ and $[Y_{0},Y_{1},\ldots,Y_{k}]$ the sampled trajectories. To capture the coupling between the controlled system and the uncontrollable agents, at each sampling instant, the movement of the uncontrollable agent, i.e., $\Delta Y_{i+1}:=Y_{i+1}-Y_{i}$, follows a distribution parameterized by the state of the controlled system, i.e.,
$\Delta Y_{i+1}\sim\mathcal{D}(X_{i}),.$ (2)
@@
Note that, distribution $\mathcal{D}(X_{i})$ is memoryless and time-invariant meaning that the behavior of the uncontrollable agent only depends on the current state of the system rather than the entire history. This assumption is valid in many applications such as the behavior of moving pedestrians. Furthermore, this distribution is unknown a priori and we can only sample data according to the distribution.
@@
Data Collections.  We assume that we have access to $\bar{K}$ independent trajectories of observations $X^{(j)}=(X_{1}^{(j)},X_{2}^{(j)},\ldots)$, $Y^{(j)}=(Y_{1}^{(j)},Y_{2}^{(j)},\ldots)$. Let $\Delta Y_{i+1}=Y_{i+1}-Y_{i}$, $\Delta Y^{(j)}=(\Delta Y_{1}^{(j)},\Delta Y_{2}^{(j)},\ldots),\forall j=1,2,\ldots,\bar{K}$ sampled from the distribution of $X^{(j)}$. Let $Z^{(j)}=(Z_{1}^{(j)},Z_{2}^{(j)},\ldots)$ where $Z_{i}^{(j)}=(X_{i}^{(j)},\Delta Y_{i}^{(j)})$. We partition these trajectories into calibration and training sets, denoted as $D_{\text{cal}}:=\{Z^{(1)},\dots,Z^{(Card(D_{cal}))}\}$ and $D_{\text{train}}:=\{Z^{(Card(D_{cal})+1)},\dots,Z^{(\bar{K})}\}$, respectively.
@@
Trajectory Predictor.  Our objective is to use the training set $D_{\text{train}}$ to design a predictor to forecast the future state of the uncontrollable agent. Note that, since $\Delta Y_{i+1}+Y_{i}=Y_{i+1}$, it suffices to predict $\Delta Y_{i+1}$ as we assume that all state information can be observed perfectly. In principle, although the $\Delta Y_{i+1}$ only depends on the current states, we still use the entire past samples to perform the prediction to improve the prediction accuracy. Therefore, the predictor is of form
$\Delta\hat{Y}_{t+1}=\Omega(Z_{(0,t]}|\theta).$ (3)
@@
Note that $\Omega$ could be implemented by any trajectory predictor such as recurrent neural networks (RNNs), long short term memory (LSTM) networks or transformers , and $\theta$ is its parameters.
IV-C Conformal Prediction for Reachable Region
@@
The predictions $\Delta\hat{Y}_{t}$, which may be generated by models with learning capabilities, are not guaranteed to be accurate. To address this, we employ conformal prediction to construct prediction regions for $\Delta\hat{Y}_{t}$ that hold with high probability.
Lemma 1
([25]). Let $V,V^{(1)},\ldots,V^{(k)}$ be $k+1$ independent and identically distributed real-valued random variables. Without loss of generality, let $V^{(1)},\ldots,V^{(k)}$ be sorted in non-decreasing order and define $V^{(k+1)}:=\infty$. For $\alpha\in(0,1)$, it holds that $\operatorname{Prob}(V\leq\bar{V})\geq 1-\alpha$ where
$\bar{V}:=V^{(r)}\text{ with }r:=\lceil(k+1)(1-\alpha)\rceil$
[…]
In the context of our approach, we aim to use conformal prediction to quantify the uncertainty in the predicted state of the uncontrollable agent.
@@
For a trajectory $Z^{(j)}=(Z_{1}^{(j)},Z_{2}^{(j)},\ldots)\in D_{cal}$ and $Z^{(j)}_{i}=(X^{(j)}_{i},\Delta Y^{(j)}_{i})$, the prediction error is then defined as
$R^{(j)}:=\sup_{Z^{(j)}_{i}\in Z^{(j)}}||\Delta Y_{i}^{(j)}-\Delta\hat{Y}_{i}^{(j)}||.$ (4)
$\Delta Y_{i}^{(j)}$ is regarded as the ground truth of $\Delta\hat{Y}_{i}^{(j)}$.
@@
Given a calibration dataset $D_{cal}\subseteq\mathbb{R}^{(n_{X},n_{Y}})$, and a failure probability tolerance $\alpha\in[0,1]$, by Lemma 1, $\forall X\in\mathbb{R}^{n_{X}}$, $t\in\mathbb{R}_{\geq 0}$, $\operatorname{Prob}\left(R^{(0)}\leq\bar{R}(\alpha)\right)\geq 1-\alpha$ holds where
$\bar{R}(\alpha)=\text{ Quantile }_{1-\alpha}\left(R^{(1)},\ldots,R^{(Card(D_{cal}))},\infty\right)$ (5)
Note that here $\operatorname{Card}(D_{\text{cal}})$ is the dataset size as in Assumption 1. Similarly, we obtain $\bar{R}=R^{(r)}$, where $r=\lceil\operatorname{Card}(D_{\text{cal}})\cdot(1-\alpha)\rceil$. As before, for meaningful prediction regions, we require $\operatorname{Card}(D_{\text{cal}})\geq\lceil(\operatorname{Card}(D_{\text{cal}})+1)(1-\alpha)\rceil$, otherwise $\bar{R}(\alpha)=\infty$.
Proof:
In Lemma 1, given a set of independent and identically distributed random variables $V^{(1)},V^{(2)},\ldots,V^{(k)}$, we can derive a bound in a sense of probability, by sorting these values and determining the corresponding quantile. In Proposition 1, we treat the nonconformity scores $R^{(1)},R^{(2)},\ldots,R^{(\text{Card}(D_{\text{cal}}))}$ as analogous to the random variables $V^{(1)},V^{(2)},\ldots,V^{(k)}$ in Lemma 1. Therefore, $R$ plays the role of $V$ in the lemma, allowing us to directly apply its result. ∎
The above result says that the state estimation error $R$ can essentially be bounded, with probability $1-\alpha$, by the worst case of conformal prediction region $\bar{R}(\alpha)$. Since the score $R^{(0)}$ represents the model’s estimation error sampled from any trajectory that might occur during the actual operation of the system, we can hence conclude that
$\operatorname{Prob}[||\Delta Y_{t+1}-\Delta\hat{Y}_{t+1}||\leq\bar{R}(\alpha)]\geq 1-\alpha$ (6)
From known measurement $X_{t}$ and $Y_{t}$, and prediction $\Delta\hat{Y}_{t+1}$, denote safe region in a probabilistic sense for $\Delta\hat{Y}_{t}$ and $\bar{R}(\alpha)$ as $\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)=[(Y_{t}+\Delta\hat{Y}_{t+1})-\bar{R}(\alpha),(Y_{t}+\Delta\hat{Y}_{t+1})+\bar{R}(\alpha)]\subseteq\mathbb{R}^{n_{Y}}$ satisfies that
$\operatorname{Prob}[Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)]\geq 1-\alpha$ (7)
Note that $\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)$ is the estimated safe region of the uncontrollable event within the next sample period.

### v3

@@ Abstract
We investigate the problem of safe control synthesis for autonomous systems operating in environments with uncontrollable agents whose dynamics are unknown but coupled with those of the controlled system. This scenario naturally arises in various applications, such as autonomous driving and human-robot collaboration, where the behavior of uncontrollable agents, such as pedestrians or human operators, cannot be directly controlled but is influenced by the actions or the states of the autonomous vehicle or robot. In this paper, we present SPARC (Safe Prediction-Based Robust Controller for Coupled Agents), a new framework designed to ensure safe control in the presence of uncontrollable agents that are coupled with the controlled system. Particularly, we introduce a joint distribution-based approach to account for the coupled dynamics of the controlled system and uncontrollable agents. Then, SPARC leverages conformal prediction to quantify the uncertainty in data-driven prediction of the agent behavior. By integrating the control barrier function (CBF), SPARC provides provable safety guarantees at a high confidence level. We illustrate our framework with a case study of an autonomous driving scenario with walking pedestrians.
@@ I INTRODUCTION
Safety is one of the major concerns in autonomous systems such as self-driving cars or industrial robots. These systems must operate in dynamic environments while avoiding unsafe behaviors. For instance, in autonomous driving, the ego vehicle must continuously monitor its surroundings to ensure the safety of itself, pedestrians, and other vehicles. To achieve safe control in real-world scenarios, it is essential to account for the presence of uncontrollable agents in the environment.
There are two major challenges in achieving safe control in such scenarios. First, the dynamics of the controlled system and the uncontrollable agents are coupled in general. In other words, the behavior of the uncontrollable agent depends on the state of the controlled system. For example, in autonomous driving, a pedestrian’s movements may change in response to the position and speed of a nearby vehicle. Second, the intentions of such agents are often unknown a priori and difficult to model precisely. As a result, data-driven approaches have emerged as a powerful tool for handling the intentions of uncontrollable agents. By leveraging large datasets collected from real-world interactions, one can develop highly accurate predictors to forecast the trajectories of uncontrollable agents for control purposes. For instance, as shown in Figure 1, in autonomous driving within pedestrian-rich environments, there are extensive datasets that capture how pedestrians react to vehicles under various conditions. However, these predictors fail to measure uncertainty in a formal and systematic manner, rendering the prediction-based control susceptible to unsafe outcomes.
(a) InD Dataset[1]
(b) Stanford Drone Dataset[2]
@@
In this paper, we propose a new framework, called SPARC (Safe Prediction-Based Robust Controller for Coupled Agents), for ensuring safe control in the presence of uncontrollable agents. Specifically, we assume the dynamics of the uncontrollable agent are unknown and coupled with the dynamics of the controlled system. To address this challenge, we leverage the theory of conformal prediction to quantify uncertainty in data-driven predictions of the uncontrollable agent’s intentions. Unlike previous approaches, which assume that the dynamics of the uncontrollable agent are independent of the controlled system, we develop a novel method to sample data and construct conformal regions based on the joint distribution over an augmented state space. This approach allows us to capture the inter-dependencies between the two agents and improve prediction accuracy. Furthermore, by incorporating the control barrier function (CBF) techniques, we can provide provable safety guarantees with a confidence bound, ensuring robust system performance. Our framework is particularly well-suited for applications such as autonomous driving in urban environments or human-robot collaboration, where uncontrollable agents operate based on their own underlying hidden decision-making processes that are influenced by the behavior of the controlled systems.
@@ II Related Works
Safety-Critical Control using CBFs. Control Barrier Functions (CBFs) are a widely used approach for ensuring safety in control systems [3, 4, 5, 6, 7]. First introduced by Ames et al. [8], the key idea behind CBFs is to encode safety constraints into a function that guarantees the system remains within a safe set during operation. However, a significant challenge of using CBFs is their reliance on an accurate dynamic model of the system. When the system includes an uncontrollable component with unknown dynamics, applying traditional CBFs can be problematic. In such cases, robust CBFs [9, 10] can be used to account for the worst-case scenario, which, however, often results in overly conservative control strategies.
@@
Conformal Prediction in Control Synthesis. Conformal prediction is an emerging technique in machine learning that provides a principle way to quantify the uncertainty in model predictions [18]. It works by generating prediction sets or confidence regions that, with a specified level of certainty, are guaranteed to contain the true outcome. This technique is model-free, meaning it can be applied to any underlying predictive model, and it offers formal statistical guarantees on the accuracy of its predictions. Recently, conformal prediction has been extensively applied to control synthesis problems; see, e.g., [19, 20, 21, 22, 23, 24]. Our work is closely related to [22] and [19], which address the problem of safe planning in unknown dynamic environments. In these works, the authors also employ conformal prediction to estimate the confidence region of the uncontrollable agent. However, a key distinction is that these approaches assume the behavior of uncontrollable agents is independent of the controlled system, an assumption that may not hold in many real-world scenarios. In contrast, our framework considers a coupled setting, where the behavior of uncontrollable agents is influenced by the state of the controlled system.
@@ III Problem Formulation
Notations: We denote $\mathbb{R},\mathbb{N}$, and $\mathbb{R}^{n}$ as the set of real numbers, natural numbers, and real vectors, respectively. Let $\beta:\mathbb{R}\rightarrow\mathbb{R}$ denote an extended class $\mathcal{K}_{\infty}$ function, i.e., a strictly increasing function with $\beta(0)=0$. For a vector $v\in\mathbb{R}^{n}$, let $|v|$, $\|v\|$, $\|v\|_{\infty}$ denote its $\ell_{1}$-norm, Euclidean norm and $\ell_{\infty}$-norm. For a finite set $D$, let $|D|$ denote its cardinality.
@@ III-A System Models
where $x(t)\in\mathbb{R}^{n}$ represents the state of the system at time $t$, and $u(t)\in\mathcal{U}$ denotes the control input with $\mathcal{U}\subseteq\mathbb{R}^{c}$ being the set of admissible control inputs. The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times c}$ represent the internal system dynamics and the input influence, respectively, and both are assumed to be locally Lipschitz continuous.
Uncontrollable Agent: We consider a scenario where the controlled system is operating in an environment with a single uncontrollable agent for the sake of simplicity. And it can be easily extended to $N$ uncontrollable agents in principle by concatenating their states and considering them as a whole. Specifically, we assume that the behavior of the uncontrollable agent is coupled with the controllable agent, i.e., the evolution of the dynamics of uncontrollable agents is conditioned on the state of the controllable agent $x(t)$. For instance, in autonomous driving scenarios involving moving pedestrians, their behavior is strongly influenced by the car’s position and velocity. Specifically, pedestrians tend to move away more quickly as the car gets closer, and they adopt more cautious walking strategies when the car approaches at higher speeds. The uncontrollable agent dynamics can be modeled as a stochastic process
$\dot{y}(t)\sim\mathbb{F}_{y}(x(t),y(t)),$ (2)
where $y(t)\in\mathbb{R}^{m}$ is the state of the uncontrollable agent, and $\mathbb{F}_{y}$ is a distribution parameterized by the states of the controlled and uncontrollable systems, which is not known a priori.
We frame safety as the 0-upper level set of a Lipschitz continuous function $h_{xy}(x,y)$, i.e., $\mathcal{S}=\{(x,y)\mid h_{xy}(x,y)\geq 0\}\subseteq\mathbb{R}^{n}\times\mathbb{R}^{m}$. And our objective is to synthesize a safety filter for reference control at each time instant such that the controlled system can ensure safety with respect to the uncontrollable agent, with a certain level of confidence.
@@
Given $h_{xy}$ defined with a safety set $\mathcal{S}$, a reference control $u_{r}$ at time $t$, we wish to obtain a safe control input $u^{*}$, that minimally interferes with the reference control while assuring safety. Formally, we aim to solve the following problem
$u^{*}$ $=\arg\min_{u\in\mathcal{U}}||u-u_{r}||^{2}$ (3)
s.t. $\text{Prob}[h_{xy}(x(t^{\prime}),y(t^{\prime}))\geq 0]\geq 1-\alpha,\;\forall t^{\prime}>t,$
[…]
Since the behavior of the uncontrollable agent is complete unknown, we will train a neural network to model its behavior and provide confidence level guarantee for safety of form: “the probability that the system becomes unsafe with the filtered input is lower than a threshold $\alpha$”.
@@ IV-A Overall Control Framework
  • At each discrete time step, referred to as a prediction update point, we use a neural network to forecast the behavior of the uncontrollable agent over the next control period $\Delta t$.
  • We apply conformal prediction techniques to expand the predicted results into a confidence region that contains the agent’s reachable states with high probability.
@@ IV-B Data Collection and Predictor Design
To collect data and to predict the behavior of the uncontrollable agent, we sample the trajectories of both the controlled system and the uncontrollable agent with time interval $\Delta t$. We denote by $[X_{0},X_{1},\ldots,X_{k}]$ and $[Y_{0},Y_{1},\ldots,Y_{k}]$ the sampled trajectories. For convenience, let $X$ denote the state space of the controllable agent which includes both position and velocity, while $Y$ represents the pedestrian’s position. To capture the coupling between the controlled system and the uncontrollable agents, at each sampling instant, the movement of the uncontrollable agent, i.e., $\Delta Y_{i}:=Y_{i+1}-Y_{i}$, follows a distribution parameterized by the state of the controlled system in (2), i.e.,
$\Delta Y_{i}\sim\mathcal{D}(X_{i},Y_{i}).$ (4)
@@
Note that distribution $\mathcal{D}(X_{i},Y_{i})$ is memoryless and time-invariant meaning that the behavior of the uncontrollable agent only depends on the current state of the system rather than the entire history. This assumption is valid in many applications such as the behavior of moving pedestrians. Furthermore, this distribution is unknown a priori and we can only sample data according to the distribution.
@@
Data Collections.  We assume that we have access to $\bar{K}$ independent trajectories of observations $X^{(j)}=(X_{1}^{(j)},X_{2}^{(j)},\ldots)$, $Y^{(j)}=(Y_{1}^{(j)},Y_{2}^{(j)},\ldots)$. Let $Z^{(j)}_{(0,t]}:=(Z_{1}^{(j)},\dots,Z_{t}^{(j)})$ where $Z_{i}^{(j)}=(X_{i}^{(j)},Y_{i}^{(j)})$.
Let $\Delta Y^{(j)}=(\Delta Y_{1}^{(j)},\Delta Y_{2}^{(j)},\ldots),\forall j=1,2,\ldots,\bar{K}$ sampled from the random distribution of $(X^{(j)},Y^{(j)})$, we denote $W^{(j)}=[(Z^{(j)}_{(0,1]},\Delta Y^{(j)}_{1}),(Z^{(j)}_{(0,2]},\Delta Y^{(j)}_{2}),\dots]$
We partition these trajectories into calibration and training sets, denoted as $D_{\text{cal}}:=\{W^{(1)},\dots,W^{(|D_{\text{cal}}|)}\}$ and $D_{\text{train}}:=\{W^{(|D_{\text{cal}}|+1)},\dots,W^{(\bar{K})}\}$, respectively.
@@
Trajectory Predictor.  Our objective is to use the training set $D_{\text{train}}$ to design a predictor to forecast the future state of the uncontrollable agent. Note that, since $\Delta Y_{i}+Y_{i}=Y_{i+1}$, it suffices to predict $\Delta Y_{i}$ as we assume that all state information can be observed perfectly. In principle, although the $\Delta Y_{i}$ only depends on the current states, we still use the entire past samples to perform the prediction to improve the prediction accuracy considering the random disturbance influence on the uncontrollable agent dynamic. Therefore, the predictor is of form
$\Delta\hat{Y}_{t}=\Omega(Z_{(0,t]}|\theta).$ (5)
@@
Note that $\Omega$ could be implemented by any trajectory predictor such as neural networks (NNs) with parameters $\theta$.
IV-C Uncertainty Qualification with Conformal Prediction
@@
The prediction $\Delta\hat{Y}_{t}$ generated by the trained model, is not guaranteed to be accurate. To address this, we employ conformal prediction to construct a prediction region for $\Delta\hat{Y}_{t}$ that holds with high probability.
@@
For a joint trajectory $Z^{(j)}_{(0,i]}=(Z_{1}^{(j)},Z_{2}^{(j)},\ldots,Z_{i}^{j})$ and $W^{(j)}_{i}=(Z^{(j)}_{(0,i]},\Delta Y^{(j)}_{i})\in D_{\text{cal}}$, the nonconformity score $R^{(j)}$ is defined as the prediction error
$R^{(j)}:=\sup_{W^{(j)}_{i}\in W^{(j)}}||\Delta Y_{i}^{(j)}-\Delta\hat{Y}_{i}^{(j)}||,$ (6)
where $\Delta Y_{i}^{(j)}$ is the ground truth of a predicted $\Delta\hat{Y}_{i}^{(j)}$.
@@
Given a calibration dataset $D_{\text{cal}}$, and a failure probability tolerance $\alpha\in[0,1]$, by [25], $\forall X\in\mathbb{R}^{n_{X}}$, $t\in\mathbb{R}_{\geq 0}$, $\operatorname{Prob}\left(R^{(0)}\leq\bar{R}(\alpha)\right)\geq 1-\alpha$ holds where
$\bar{R}(\alpha)=\text{ Quantile }_{1-\alpha}\left(R^{(1)},\ldots,R^{(|D_{\text{cal}}|)},\infty\right)$ (7)
Note that here $\bar{R}(\alpha)\in\mathbb{R}^{m}$ is a vector for each dimension and $|D_{\text{cal}}|$ is the calibration dataset size. Similarly, we obtain $\bar{R}=R^{(r)}$, where $r=\lceil|D_{\text{cal}}|\cdot(1-\alpha)\rceil$. As before, for meaningful prediction regions, we require $|D_{\text{cal}}|\geq\lceil(|D_{\text{cal}}|+1)(1-\alpha)\rceil$, otherwise $\bar{R}(\alpha)=\infty$.
Conformal prediction implies the nonconformity score $R$ can essentially be bounded, with probability $1-\alpha$, within the conformal region $\bar{R}(\alpha)$. Given that the test and calibration data are exchangeable, we can hence conclude that for each coordinate direction $i=1,...,m,$
$\operatorname{Prob}\left[\left(\frac{|y_{i}-(Y_{t}+\Delta\hat{Y}_{t})_{i}|}{\bar{R}_{i}(\alpha)}\right)\leq 1\right]\geq 1-\alpha$ (8)
From known measurement $X_{t}$ and $Y_{t}$, and prediction $\Delta\hat{Y}_{t}$, we denote the conformal region for $Y_{t+1}$ in a probabilistic sense as
$\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)=\left\{y\in\mathbb{R}^{m}\mid\max_{i}\left(\frac{|y_{i}-(Y_{t}+\Delta\hat{Y}_{t})_{i}|}{\bar{R}_{i}(\alpha)}\right)\leq 1\right\}$ (9)
This set represents a weighted $\ell_{\infty}$-norm ball centered at $(Y_{t}+\Delta\hat{Y}_{t})$, where $\bar{R}(\alpha)$ determines the scaling in each coordinate direction. It satisfies that
$\operatorname{Prob}[Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)]\geq 1-\alpha.$ (10)
Note that $\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)$ is the estimated conformal prediction region of the uncontrollable agent in the next sampling step.